# A General Framework for the Spontaneous Scalarization of Regular Black Holes

### v1

@@ I Introduction
Black holes (BHs) are a core prediction of the theory of general relativity (GR). In electrovacuum, the Reissner-Nordström (RN) and Kerr-Newman solutions uniquely describe these objects and yet host singularities [1, 2]. This uncomfortable fact has led to investigations of how “hair” may be introduced into the action to regularize them (see Ref. [3] for a review). One such path involves nonlinear electrodynamics (NLED) [4, 5, 6]: although Coulombic repulsion is not strong enough to prevent gravitational collapse, if the electromagnetic sector is augmented in the high-energy limit the attractive nature of gravity can be counteracted at some characteristic length-scale in such a way that collapse halts before a singularity forms for arbitrarily small charge.
@@
In this work, we study the spontaneous scalarization that arises from the coupling between a regular BH solution supported by NLED and a nonminimally coupled scalar field. The method is general and works for an arbitrary (static) seed, and therefore generalizes studies that have considered a similar problem but in the restricted contexts of the RN [19], power-Maxwell [12], Bardeen [20], or Euler-Heisenberg [21] BHs. After providing a general recipe and discussing numerical techniques, the method is demonstrated for a simple case with a regular BH (namely the Balart-Vagenas solution [22]). We explore some astrophysical manifestations for the new solutions found here, for example as concerns shadows and quasi-normal modes (QNMs).
@@ II.1 Inverse problem and the P-dual formalism
Note that, the Maxwell case is recovered when $\mathcal{H}_{P}=1$, as expected.
@@
Given a choice of $N$, one need only solve equations (10) and (12) for $F(r)$ and $\mathcal{L}_{\rm EM}$ (utilising 7), respectively, to find an exact theory around the metric. A specific example of a well-motivated background is considered in Sec. IV, though we now turn to the main question posed in this work: how does a general, regular BH respond to a non-minimal scalar coupling?
@@ III Spontaneous scalarization
At this point, a few remarks are in order. First, note that when $\mathcal{H}_{P}=1$, the set of equations (24)–(27) reduces to the standard Maxwell case reported in [19]. Secondly, Eqs. (24)–(27) are generally valid for any $\mathcal{H}(P)$, provided the assumptions of staticity and the absence of magnetic monopoles hold. In this regard, for any given model $\mathcal{H}(P)$, one can insert it into the above expressions. Next, note that the equations depend only on the radial coordinate. In fact, from (5) we obtain
@@ III.1 Domain of existence
Before attempting to solve the full system (25)–(27) together with the scalar equation (18), we examine the domain of existence. This set corresponds to the valid range of metric and theory parameters such that a scalarized branch of solutions exists; identifying this set simplifies the subsequent numerical analysis (see Sec. III.2).
In general, a scalar field perturbation over a static and spherically-symmetric background can be written
@@
If $\mu_{\rm eff}<0$, the tachyonic instability can occur as described in Sec. III. The solution of (31) together with Dirichlet boundary conditions to fix $\omega_{\ell m}$ defines the lower bound – the so-called existence line – for scalarized solutions: a given set of parameters sets the minimal, necessary conditions for exponentially-growing modes. In the simple case we are considering, we need only consider the fundamental $\ell=0$ eigenfunction ($U_{0}$), as this defines the strongest modes, so that the task of obtaining the existence line reduces to identifying the zeros of $U_{0}$ at infinity ($r\to\infty$).
@@ III.2 Numerical methods
The system (25)–(27) cannot be solved analytically in general; therefore, we employ the shooting method (see, for instance, [19, 12] and references therein for applications). This method consists of an iterative procedure in which initial trial values for the unknowns are adjusted until the boundary conditions are satisfied. In our model, the functions $m,\delta,\phi$} must be determined for any $r\in(r_{H},\infty)$. Since we require asymptotically flat solutions, the boundary conditions at infinity are $m\to M$, $\delta\to 0$, and $\phi\to 0$. However, the exact values of $\delta$ and $\phi$ at the horizon $r_{H}$ remain undetermined. To reduce the number of free parameters, we expand the functions near the horizon as
@@
$\int\limits_{r_{H}}^{\infty}drr^{2}e^{-\delta-\alpha\phi^{2}}\left[\left(3-\frac{2r_{H}}{r}\right)\mathcal{H}-4P\mathcal{H}_{P}\left(1-\frac{r_{H}}{r}\right)\right]$ (38)
$\hskip 40.0pt=\int\limits_{r_{H}}^{\infty}dr\frac{e^{-\delta}r^{2}\phi^{\prime}}{2}\left[1+\frac{2r_{H}}{r}\left(\frac{m}{r}-1\right)\right]$
Both sides of expression (38) are checked for each numerical solution, and if the mismatch lies below a tolerance (set as $10^{-6}$ in appropriate, dimensionless units) the result is considered acceptable.
@@ IV Worked example: Balart-Vagenas black hole
The domain of existence is obtained through numerical iteration by fixing $\alpha$ and $Q$, and varying $r_{H}$. For each $r_{H}$, the equations of motion are solved, and the initial guess for $\phi_{0}$ is taken from a neighboring solution. Each $\alpha$-branch starts at the existence line (obtained by integrating (31)) and ends at the critical set, defined by a vanishing horizon area. The scalarized solutions exhibit a virial mismatch on the order of $10^{-6}$. The domain of existence is shown in Fig. 2.
It is worth noting that there exists a region of non-uniqueness within the domain of existence where scalar-free and scalarized BHs coexist, as defined by the values of $q=Q/m_{\infty}$ for which the metric $N$ possesses real roots. In this region, the scalarized solutions are entropically preferred, as they maximize the entropy (or, equivalently, the horizon area $A_{H}$), as shown in Fig. 3 for the specific values indicated in the legend. Moreover, the entropically preferred solutions correspond to those for which the coupling parameter $|\alpha|$ increases.
@@ VI Conclusions
The framework presented here is broadly applicable to any static seed geometry supported by nonlinear electrodynamics and readily extends to other coupling choices $f(\phi)$, provided the Bekenstein identities and the relevant energy conditions hold. Natural next steps include: (i) stability analyses beyond the virial test (time-domain evolutions and full mode spectra), (ii) extensions to dyonic and rotating spacetimes, (iii) dynamical formation and transition between branches, (iv) refined constraints from shadows, ringdowns, and strong-field lensing, and (v) other models inspired in alternative approaches [42, 43]. These avenues will sharpen the viability of scalarized, nonlinear electrodynamics-supported BHs as signatures of beyond–GR physics in the strong-field regime.
@@ Acknowledgements.
E. C. and P.B acknowledges financial support from the Generalitat Valenciana through PROMETEO PROJECT CIPROM/2022/13. E. C. is funded by the Beatriz Galindo contract BG23/00163 (Spain). M. C. H. thanks ICTP (Italy) where this work was initiated. AGS acknowledges funding from the European Union’s Horizon MSCA-2022 research and innovation programme “EinsteinWaves” under grant agreement No. 101131233.

### v2

@@ I Introduction
Black holes (BHs) are a core prediction of the theory of general relativity (GR). In electrovacuum, the Reissner-Nordström (RN) and Kerr-Newman solutions uniquely describe these objects and yet host singularities [1, 2]. This uncomfortable fact has led to investigations of how “hair” may be introduced into the action to regularize them (see Ref. [3] for a review). One such path involves nonlinear electrodynamics (NLED) [4, 5, 6, 7]: although Coulombic repulsion is not strong enough to prevent gravitational collapse, if the electromagnetic sector is augmented in the high-energy limit the attractive nature of gravity can be counteracted at some characteristic length-scale in such a way that collapse halts before a singularity forms for arbitrarily small charge.
@@
In this work, we study the spontaneous scalarization that arises from the coupling between a regular BH solution supported by NLED and a nonminimally coupled scalar field. The method is general and works for an arbitrary (static) seed, and therefore generalizes studies that have considered a similar problem but in the restricted contexts of the RN [20], power-Maxwell [13], Bardeen [21, 22], or Euler-Heisenberg [23] BHs. After providing a general recipe and discussing numerical techniques, the method is demonstrated for a simple case with a regular BH (namely the Balart-Vagenas solution [24]). We explore some astrophysical manifestations for the new solutions found here, for example as concerns shadows and quasi-normal modes (QNMs).
@@ II.1 Inverse problem and the P-dual formalism
${\nabla_{\mu}\left(\frac{d\mathcal{L}_{\text{EM}}}{dF}\,F^{\mu\nu}\right)=0,}$ (4)
from which it follows that
@@
Note that the Maxwell case is recovered when $\mathcal{H}_{P}=1$, as expected.
@@
Given a choice of $N$, one need only solve equations (11) and (13) for $F(r)$ and $\mathcal{L}_{\rm EM}$ (utilising 8), respectively, to find an exact theory around the metric. A specific example of a well-motivated background is considered in Sec. IV, though we now turn to the main question posed in this work: how does a general, regular BH respond to a non-minimal scalar coupling?
@@ III Spontaneous scalarization
At this point, a few remarks are in order. First, note that when $\mathcal{H}_{P}=1$, the set of equations (25)–(28) reduces to the standard Maxwell case reported in Herdeiro et al. [20]. Secondly, Eqs. (25)–(28) are generally valid for any $\mathcal{H}(P)$, provided the assumptions of staticity and the absence of magnetic monopoles hold. In this regard, for any given model $\mathcal{H}(P)$, one can insert it into the above expressions. Next, note that the equations depend only on the radial coordinate. In fact, from (6) we obtain
@@ III.1 Domain of existence
Before attempting to solve the full system (26)–(28) together with the scalar equation (19), we examine the domain of existence. This set corresponds to the valid range of metric and theory parameters such that a scalarized branch of solutions exists; identifying this set simplifies the subsequent numerical analysis (see Sec. III.2). In general, a scalar field perturbation over a static and spherically-symmetric background can be written
@@
$\frac{e^{2\delta}\omega^{2}_{lm}}{N}U_{\ell}=\frac{e^{\delta}}{r^{2}}\frac{d}{dr}\left(\frac{r^{2}N}{e^{\delta}}\frac{dU_{\ell}}{dr}\right)-\left[\frac{\ell(\ell+1)}{r^{2}}+\mu_{\rm eff}^{2}\right]U_{\ell}.$ (32)
Now, a necessary condition for the tachyonic instability to occur is that frequency becomes imaginary. Thus, the boundary between positive and negative values of $\omega^{2}_{\ell m}$ is defined by the condition $\omega^{2}_{\ell m}=0$, which leads to
@@
If $\mu_{\rm eff}<0$, the tachyonic instability can occur as described in Sec. III. The solution of equation (33) together with Dirichlet boundary conditions to fix $\omega_{\ell m}$ defines the lower bound – the so-called existence line – for scalarized solutions: a given set of parameters sets the minimal, necessary conditions for exponentially-growing modes. In the simple case we are considering, we need only consider the fundamental $\ell=0$ eigenfunction ($U_{0}$), as this defines the strongest modes, so that the task of obtaining the existence line reduces to identifying the zeros of $U_{0}$ at infinity ($r\to\infty$).
@@ III.2 Numerical methods
The system (26)–(28) cannot be solved analytically in general; therefore, we employ the shooting method (see, for instance, Refs. [20, 13] and references therein for applications). This method consists of an iterative procedure in which initial trial values for the unknowns are adjusted until the boundary conditions are satisfied. In our model, the functions $m,\delta,\phi$} must be determined for any $r\in(r_{H},\infty)$. Since we require asymptotically flat solutions, the boundary conditions at infinity are $m\to M$, $\delta\to 0$, and $\phi\to 0$. However, the exact values of $\delta$ and $\phi$ at the horizon $r_{H}$ remain undetermined. To reduce the number of free parameters, we expand the functions near the horizon as
@@
$\int\limits_{r_{H}}^{\infty}drr^{2}e^{-\delta-\alpha\phi^{2}}\left[\left(3-\frac{2r_{H}}{r}\right)\mathcal{H}-4P\mathcal{H}_{P}\left(1-\frac{r_{H}}{r}\right)\right]$
$\hskip 40.0pt=\int\limits_{r_{H}}^{\infty}dr\frac{e^{-\delta}r^{2}\phi^{\prime}}{2}\left[1+\frac{2r_{H}}{r}\left(\frac{m}{r}-1\right)\right]$ (40)
Both sides of expression (III.2) are checked for each numerical solution, and if the mismatch lies below a tolerance (set as $10^{-6}$ in appropriate, dimensionless units) the result is considered acceptable.
@@ IV Worked example: Balart-Vagenas black hole
The domain of existence is obtained through numerical iteration by fixing $\alpha$ and $Q$, and varying $r_{H}$. For each $r_{H}$, the equations of motion are solved, and the initial guess for $\phi_{0}$ is taken from a neighboring solution. Each $\alpha$-branch starts at the existence line (obtained by integrating (33)) and ends at the critical set, defined by a vanishing horizon area. The scalarized solutions exhibit a virial mismatch on the order of $10^{-6}$. The domain of existence is shown in Fig. 2. At this point it is important to highlight that we have explicitly checked the numerical convergence of our solutions. In particular, we repeated the integration by using different resolutions (varying the radial step size) and verified that the deviation from the virial identity systematically decreases as the step size is reduced as shown in Fig. 3. For instance, choosing $\Delta r\leq 0.006$ (again in dimensionless units) shows that the virial identity is numerically satisfied within our specified tolerance of $10^{-6}$ for any values of $\alpha\leq 50$. Moreover, the monotonic reduction of the mismatch as a function of decreasing $\Delta r$ shows the scheme converges with increasing resolution, as expected.
Figure 3: Virial identity as a function of the numerical step size. The results show that the virial identity converges to zero in the limit of vanishing step size.
It is worth noting that there exists a region of non-uniqueness within the domain of existence where scalar-free and scalarized BHs coexist, as defined by the values of $q=Q/m_{\infty}$ for which the metric $N$ possesses real roots. In this region, the scalarized solutions are entropically preferred, as they maximize the entropy (or, equivalently, the horizon area $A_{H}$), as shown in Fig. 4 for the specific values indicated in the legend. It is clear that the entropically-preferred solutions correspond to those for which the coupling parameter $|\alpha|$ increases.
@@ VI Conclusions
The framework presented here is broadly applicable to any static seed geometry supported by nonlinear electrodynamics and readily extends to other coupling choices $f(\phi)$, provided the Bekenstein identities and the relevant energy conditions hold. Natural next steps include: (i) stability analyses beyond the virial test (time-domain evolutions and full mode spectra), (ii) extensions to dyonic and rotating spacetimes, (iii) dynamical formation and transition between branches, (iv) refined constraints from shadows, ringdowns, and strong-field lensing, and (v) other models inspired in alternative approaches [44, 45]. These avenues will sharpen the viability of scalarized, nonlinear electrodynamics-supported BHs as signatures of beyond–GR physics in the strong-field regime (including, perhaps, for those with nontrivial topology [46, 47] or non-Minkowski cores [48]).
@@ Acknowledgements.
We thank the anonymous referee for their helpful comments. E. C. and P.B acknowledge financial support from the Generalitat Valenciana through PROMETEO PROJECT CIPROM/2022/13. E. C. is funded by the Beatriz Galindo contract BG23/00163 (Spain). M. C. H. thanks ICTP (Italy) where this work was initiated. AGS acknowledges funding from the European Union’s Horizon MSCA-2022 research and innovation programme “EinsteinWaves” under grant agreement No. 101131233.